# Learnability in Online Kernel Selection with Memory Constraint via Data-dependent Regret Analysis

### v1

@@ 4.2 Smooth Loss Functions
and $\lambda_{i}=\frac{2U}{G_{1}\sqrt{B}}$. Let $\mathcal{R}$ and $\ell$ satisfy Lemma 2. With probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies
$\forall\kappa_{i}\in\mathcal{K},~{}\forall f$ $\in\mathbb{H}_{i},\quad\mathrm{Reg}(f)\leq$
$\frac{24UG_{2}L_{T}(f)}{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}+UG_{1}\sqrt{B}+\frac{100U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+10U(1-\gamma)^{-\frac{3}{2}}\sqrt{G_{2}G_{1}\ln\frac{1}{\delta}}\sqrt{L_{T}(f)+\frac{UG_{1}\sqrt{B}}{4}},$
where $\gamma={6UG_{2}}/{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}$.
In the case of $B\ll T$, the dominant term in the upper bound is $O\left(\frac{L_{T}(f)}{\sqrt{B}}\right)$, Previous work proved a $\Omega(T/B)$ lower bound [19]. Thus our result proves a new trade-off between memory constraint and regret bound which gives an answer to Q1. If $\min_{f\in\mathbb{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$, then a $\Theta(\ln{T})$ memory budget is enough to achieve a sub-linear regret which answers Q2. According to Definition 2, we also conclude that $\mathbb{H}_{\kappa^{\ast}}$ is online learnable. The penalty term for not knowing the optimal kernel is $O(U\sqrt{L_{T}(f)\ln{K}})$ which only depends on $O(\sqrt{\ln{K}})$. In this case, online kernel selection is not much harder than online kernel learning.
Theorem 3 can also recover some state-of-the-art results. Let $B=\Theta(T)$. We obtain a $O(U\sqrt{T\ln{K}})$ regret bound which is same with the regret bound in [1, 2, 10]. If $K=1$ and $B=\Theta(L_{T}(f))$, then we obtain a $O(U\sqrt{L_{T}(f)})$ regret bound matching the regret bound in [19]. Next we prove that M-OMD-S is optimal.
@@
Our lower bound improves the previous $\Omega(T/B)$ [19]. Next we show an algorithm-dependent bound. Let
@@
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}|$. It is natural that if $(f_{t,i}(2)-f)$ is close to $(f_{t,i}-f)$, then $\xi_{i}\ll 4$. The restart technique sets $f_{t,i}(2)=0$ which implies $\Lambda_{i}\leq U^{2}|\mathcal{J}|$.
@@
Let the conditions in Theorem 3 be satisfied and $U\leq\frac{B-\frac{4}{3}\ln\frac{1}{\delta}}{32G_{2}}$. For a fixed $i\in[K]$, if $\xi_{i}<\frac{1}{|\mathcal{J}|}$, then with probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies,
$\forall f\in\mathbb{H}_{i},\mathrm{Reg}(f)\leq$ $\frac{32UG_{2}L_{T}(f)}{B-\frac{4}{3}\ln\frac{1}{\delta}}+\frac{144U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+12U(1-\gamma)^{-\frac{3}{2}}\sqrt{G_{2}G_{1}L_{T}(f)\ln\frac{1}{\delta}},$
@@
The dominate term in Theorem 5 is $O(\frac{UG_{2}L_{T}(f)}{B})$, while the dominate term in Theorem 3 is $O(\frac{UG_{2}L_{T}(f)}{\sqrt{B}})$. The restart technique can only enjoys a regret bound given in Theorem 3. Thus removing a half of examples can be better than restart.
@@ 5 Experiments
    We will verify the data-dependent complexities are smaller than the worst-case complexity, i.e., $\min_{\kappa_{i}}\mathcal{A}_{T,\kappa_{i}}\ll T$ and $\min_{f\in\mathbb{H}_{i},i\in[K]}L_{T}(f)\ll T$.
@@ 5.1 Experimental Setting
We implement M-OMD-S with the logistic loss function. Thus both M-OMD-H and M-OMD-S are suitable for online binary classification tasks. We adopt the Gaussian kernel $\kappa({\bm{x}},{\bm{v}})=\exp(-\frac{\|{\bm{x}}-{\bm{v}}\|^{2}_{2}}{2\sigma^{2}})$, and choose 8 classification datasets from UCI machine learning repository  [^2] http://archive.ics.uci.edu/ml/datasets.php. The information of the datasets is shown in Table 1. All algorithms are implemented in R on a Windows machine with 2.8 GHz Core(TM) i7-1165G7 CPU. We execute each experiment 10 times with random permutation of all datasets.
@@
We compare with two baseline algorithms.
  • Raker: online multi-kernel learning via random features [10].
  • LKMBooks: OKS with memory constraint via random adding examples [4].
Our algorithms improve the regret bounds of Raker and LKMBooks, and we expect that our algorithms also perform better. Although there are other online kernel selection algorithms, such as OKS [26], OMKC [27], OMKR [1], we do not compare with these algorithms, since they do not limit the memory budget.
We select $K=5$ Gaussian kernel function by setting $\sigma=2^{[-2:2:6]}$. For the baseline algorithms, we tune the hyper-parameters following original papers. To be specific, for LKMBooks, there are three hyper-parameters $\nu\in(0,1)$, $\lambda$ and $\eta$. Both $\lambda$ and $\eta$ are learning rates. We reset $\lambda=10\frac{\sqrt{(1+\nu)B}}{\sqrt{1-\nu}T}$ for improving performance and tune $\nu\in(\frac{1}{2},\frac{1}{3},\frac{1}{4})$. For Raker, there is a learning rate $\eta$ and a regularization parameter $\lambda$. We tune $\eta=\frac{10^{-3:1:3}}{\sqrt{T}}$ and $\lambda\in\{0.05,0.005,0.0005\}$. For M-OMD-H and M-OMD-S, we can set an aggressive value of $U=\sqrt{B}$ following Theorem 2 and Theorem 5. We tune the learning rate $\lambda_{i}=c\frac{U}{\sqrt{B}}$ where $c\in\{2,1,0.5\}$. The value of $\eta_{t}$ follows (10). The value of $\gamma_{t}$ follows (12) and (17). For M-OMD-H, we set $M=10$.
Let $\mathrm{MR}=\frac{1}{T}\sum^{T}_{t=1}\mathbb{I}_{\hat{y}_{t}\neq y_{t}}$ be the mistake rate. Table 2 and Table 3 summary the experimental results including the average MR (AMR) of 10 random running, memory budget ($B$ or $D$), $\mathcal{A}_{T,\kappa^{\ast}}=\min_{\kappa_{i}}\mathcal{A}_{T,\kappa_{i}}$ and $L^{\ast}=\min_{f\in\mathbb{H}_{i},i\in[k]}L_{T}(f)$. $D$ is the number of random features adopted by Raker and plays a similar role with $B$. For completeness, we also report the average running time which is not a critical measurement.
@@ 5.2.1 Hinge Loss Function
Table 2 shows the results. Note that M-OMD-H keeps $5$ buffers, i.e., $S_{i},i=1,\ldots,5$, $|S_{i}|\leq B$ and LKMBooks only keeps a buffer $S$, $|S|\leq B$. Thus Raker, LKMBooks and M-OMD-H with $B=100$ enjoy a same memory budget.
We first compare the AMR of all algorithms. As a whole, M-OMD-H with $B=400$ performs best on most of datasets, while M-OMD-H with $B=100$ performs worst. The reason is that M-OMD-H uniformly allocates the memory budget over all kernels which induces a $O(\sqrt{K})$ factor on the regret bound. If we increase the memory by $K$ times, then the $O(\sqrt{K})$ factor can be removed. Thus M-OMD-H with $B=400$ performs better. The experimental results do not completely coincide with G2. It is left for future work to study whether we can avoid allocating the memory budget.
Next we analyze the value of $\mathcal{A}_{T,\kappa^{\ast}}$. Denote by $\hat{\mathcal{A}}_{T,i}:=\sum_{\tau\leq T}\left\|\nabla_{\tau,i}-\hat{\nabla}_{\tau,i}\right\|^{2}_{\mathcal{H}_{i}}\cdot\mathbb{I}_{\nabla_{\tau,i}\neq 0}$. Computing $\mathcal{A}_{T,\kappa^{\ast}}$ requires time in $O(T^{2})$. For simplicity, we use $\min_{i\in[K]}\hat{\mathcal{A}}_{T,i}$ as a proxy of $\mathcal{A}_{T,\kappa^{\ast}}$. In fact, our regret bound depends on $\hat{\mathcal{A}}_{T,i}$ which satisfies $\hat{\mathcal{A}}_{T,i}=\tilde{O}(\mathcal{A}_{T,\kappa_{i}})$. To obtain a precise estimation of $\mathcal{A}_{T,\kappa_{i}}$, we again run M-OMD-H with $M=30$ and $B=400$.
We find that $\mathcal{A}_{T,\kappa^{\ast}}<T$ on all datasets. Especially, on the mushrooms, phishing, w8a datasets, we have $\mathcal{A}_{T,\kappa^{\ast}}\ll T$. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
@@
Raker [10] 11.97 $\pm$ 0.93 400 1.57 - 9.13 $\pm$ 0.28 400 2.06 -
LKMBooks [4] 6.04 $\pm$ 0.62 400 1.64 - 12.81 $\pm$ 0.54 400 2.21 -
@@
Raker [10] 32.14 $\pm$ 0.43 400 3.25 - 23.47 $\pm$ 0.23 400 9.45 -
LKMBooks [4] 23.66 $\pm$ 0.51 400 1.31 - 21.29 $\pm$ 1.36 400 10.15 -
@@
Raker [10] 2.98 $\pm$ 0.00 400 11.62 - 29.62 $\pm$ 0.46 400 8.88 -
LKMBooks [4] 2.97 $\pm$ 0.00 400 23.52 - 29.37 $\pm$ 0.87 400 5.47 -
@@
Raker [10] 9.49 $\pm$ 0.08 400 23.54 - 11.90 $\pm$ 0.31 400 53.32 -
LKMBooks [4] 9.58 $\pm$ 0.01 400 18.85 - 12.57 $\pm$ 0.26 400 19.17 -
@@ 5.2.2 Logistic Loss Function
Table 3 summaries the results. As a whole, M-OMD-S performs best on most of datasets. On the phishing, SUSY and cod-rna datasets, Raker performs better than M-OMD-S. We find that the current regret analysis of Raker is not tight, and can be improved using Assumption 3. Thus it is reasonable that Raker performs similar with M-OMD-S. Besides, M-OMD-S performs better than LKMBooks on all datasets. The experimental results coincide with the theoretical analysis that is, M-OMD-S enjoys a better regret bound than LKMBooks. The experimental results verify G2.
Next we analyze the value of $L^{\ast}$. For simplicity, we use the cumulative losses of M-OMD-S, $\hat{L}_{1:T}$, as a proxy of $L^{\ast}$. Note that $L^{\ast}\leq\hat{L}_{1:T}$. Table 3 shows that $L^{\ast}<T$ on all datasets. Besides, on the mushrooms, phishing and w8a datasets, $L^{\ast}\ll T$. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
@@
Raker [10] 11.47 $\pm$ 0.75 400 1.64 - 9.07 $\pm$ 0.44 400 2.05 -
LKMBooks [4] 10.01 $\pm$ 1.44 400 2.12 - 15.86 $\pm$ 1.28 400 1.94 -
@@
Raker [10] 30.74 $\pm$ 1.11 400 3.56 - 23.13 $\pm$ 0.17 400 9.90 -
LKMBooks [4] 25.72 $\pm$ 0.64 400 1.30 - 22.38 $\pm$ 1.29 400 10.98 -
@@
Raker [10] 2.97 $\pm$ 0.00 400 12.06 - 28.62 $\pm$ 0.44 400 9.02 -
LKMBooks [4] 2.98 $\pm$ 0.01 400 6.10 - 30.72 $\pm$ 1.33 400 6.45 0
@@
Raker [10] 9.53 $\pm$ 0.03 400 24.94 - 12.27 $\pm$ 0.57 400 52.24 -
LKMBooks [4] 9.57 $\pm$ 0.00 400 16.95 - 13.91 $\pm$ 1.31 400 25.55 -

### v2

@@ 5.2 Smooth Loss Functions
and $\lambda_{i}=\frac{2U}{G_{1}\sqrt{B}}$ for all $i\in[K]$. With probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies
$\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i},$ $\quad\mathrm{Reg}(f)\leq$
$\frac{24UG_{2}L_{T}(f)}{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}+UG_{1}\sqrt{B}+\frac{100U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+\frac{10U}{(1-\gamma)^{\frac{3}{2}}}\sqrt{G_{2}G_{1}\ln\frac{1}{\delta}}\sqrt{L_{T}(f)+\frac{UG_{1}\sqrt{B}}{4}},$
where $\gamma=\frac{6UG_{2}}{\sqrt{B-\frac{4}{3}\ln\frac{1}{\delta}}}$.
In the case of $B\ll T$, the dominant term in the upper bound is $O\left(\frac{L_{T}(f)}{\sqrt{B}}\right)$. Previous work proved a $\Omega\left(\frac{T}{B}\right)$ lower bound [Zhang2013Online]. Thus our result proves a new trade-off between memory constraint and regret which gives an answer to Q1. If $\min_{f\in\mathbb{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$, then a $\Theta(\ln{T})$ memory budget is enough to achieve a sub-linear regret which answers Q2. By Definition 2, we also conclude that $\mathbb{H}_{\kappa^{\ast}}$ is online learnable. The penalty term for not knowing the optimal kernel is $O\left(U\sqrt{L_{T}(f)\ln{K}}\right)$ which only depends on $O\left(\sqrt{\ln{K}}\right)$. In this case, online kernel selection is not much harder than online kernel learning.
Theorem 3 can also recover some state-of-the-art results. Let $B=\Theta(T)$. We obtain a $O\left(U\sqrt{T\ln{K}}\right)$ regret bound which is same with the regret bound in [Sahoo2014Online, Foster2017Parameter, Shen2019Random]. If $K=1$ and $B=\Theta(L_{T}(f))$, then we obtain a $O\left(U\sqrt{L_{T}(f)}\right)$ regret bound, matching the regret bound in [Zhang2013Online]. Next we prove that M-OMD-S is optimal.
@@
Our lower bound improves the previous lower bound $\Omega\left(\frac{T}{B}\right)$ [Zhang2013Online]. Next we show an algorithm-dependent upper bound. Let
@@
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}|$. It is natural that if $(f_{t,i}(2)-f)$ is close to $(f_{t,i}-f)$, then $\xi_{i}\ll 4$. The restart technique sets $f_{t,i}(2)=0$ implying $\Lambda_{i}\leq U^{2}|\mathcal{J}|$.
@@
Suppose the conditions in Theorem 3 are satisfied. Let $\delta\in(0,1)$ and $U\leq\frac{B-\frac{4}{3}\ln\frac{1}{\delta}}{32G_{2}}$. If $\xi_{i}<\frac{1}{|\mathcal{J}|}$ for all $i\in[K]$, then with probability at least $1-\frac{3T}{\alpha\mathcal{R}}\Theta(\lceil\ln{T}\rceil)\delta$, the regret of M-OMD-S satisfies,
$\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i},\quad\mathrm{Reg}(f)\leq\frac{32UG_{2}L_{T}(f)}{B-\frac{4}{3}\ln\frac{1}{\delta}}+\frac{144U^{2}G_{2}G_{1}\ln\frac{1}{\delta}}{(1-\gamma)^{2}}+\frac{12U}{(1-\gamma)^{\frac{3}{2}}}\sqrt{G_{2}G_{1}L_{T}(f)\ln\frac{1}{\delta}},$
@@
The dominate term in Theorem 5 is $O\left(\frac{UG_{2}L_{T}(f)}{B}\right)$, while the dominate term in Theorem 3 is $O\left(\frac{UG_{2}L_{T}(f)}{\sqrt{B}}\right)$. The restart technique only enjoys the regret bound given in Theorem 3. Thus removing a half of examples can be better than the restart technique.
@@ 6 Experiments
    We aim to verify the two data-dependent complexities are smaller than the worst-case complexity $T$, that is, $\min_{\kappa_{i},i\in[K]}\mathcal{A}_{T,\kappa_{i}}\ll T$ and $\min_{f\in\mathbb{H}_{i},i\in[K]}L_{T}(f)\ll T$.
@@ 6.1 Experimental Setting
We implement M-OMD-S with the logistic loss function. Both M-OMD-H and M-OMD-S are suitable for online binary classification tasks. We adopt the Gaussian kernel $\kappa({\bm{x}},{\bm{v}})=\exp(-\frac{\|{\bm{x}}-{\bm{v}}\|^{2}_{2}}{2\sigma^{2}})$, and choose eight classification datasets. The information of the datasets is shown in Table 1. The w8a, the ijcnn1, and the cod-rna datasets are download from the LIBSVM website  [^2] https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/, Jul. 2024.. The other datasets are downloaded from UCI machine learning repository  [^3] http://archive.ics.uci.edu/ml/datasets.php, Jul. 2024.. All algorithms are implemented in R on a Windows machine with 2.8 GHz Core(TM) i7-1165G7 CPU. We execute each experiment ten times with random permutation of all datasets.
@@
We compare our algorithms with two baseline algorithms.
  • LKMBooks: OKS with memory constraint via random adding examples [Li2022Worst].
  • Raker: online multi-kernel learning using random features [Shen2019Random].
Our algorithms improve the regret bounds of Raker and LKMBooks, and we expect that our algorithms also perform better. Although there are other online kernel selection algorithm, such as OKS [Yang2012Online], and online multi-kernel learning algorithms, such as OMKR [Sahoo2014Online] and OMKC [Hoi2013Online], we do not compare with these algorithms, since they do not limit the memory budget.
We select $K=5$ Gaussian kernel function by setting $\sigma=2^{[-2:2:6]}$. For the baseline algorithms, we tune the hyper-parameters following original papers. To be specific, for LKMBooks, there are three hyper-parameters $\nu\in(0,1)$, $\lambda$ and $\eta$. Both $\lambda$ and $\eta$ are learning rates. We reset $\lambda=\frac{10}{\sqrt{1-\nu}T}\sqrt{(1+\nu)B}$ for improving performance and tune $\nu\in(\frac{1}{2},\frac{1}{3},\frac{1}{4})$. For Raker, there is a learning rate $\eta$ and a regularization parameter $\lambda$. We tune $\eta=\frac{10^{-3:1:3}}{\sqrt{T}}$ and $\lambda\in\{0.05,0.005,0.0005\}$. For M-OMD-H and M-OMD-S, we can set an aggressive value of $U=\sqrt{B}$ following Theorem 2 and Theorem 5. We tune the learning rate $\lambda_{i}=c\frac{U}{\sqrt{B}}$ where $c\in\{2,1,0.5\}$. The value of $\eta_{t}$ follows (10). The value of $\gamma_{t}$ follows (12) and (17). For M-OMD-H, we set $M=10$.
We always set $D=400$ in Raker where $D$ is the number of random features adopted by Raker and plays a similar role with $B$. For LKMBooks and M-OMD-S, we always set the budget $B=400$. Note that M-OMD-H uniformly allocates the memory budget over $K$ hypotheses and the reservoir. Thus we separately set $B=100$ and $B=400$ for M-OMD-H, denoted as M-OMD-H $(B=100)$ and M-OMD-H $(B=400)$. Raker, LKMBooks, M-OMD-S, and M-OMD-H $(B=100)$ use a same memory budget, while M-OMD-H $(B=400)$ uses a larger memory budget.
@@ 6.2.1 Hinge Loss Function
We first compare the AMR of all algorithms. Table 2 shows the results. As a whole, M-OMD-H $(B=400)$ performs best on most of datasets, while M-OMD-H $(B=100)$ performs worst. The reason is that M-OMD-H uniformly allocates the memory budget over all kernels thereby inducing a $O\left(\sqrt{K}\right)$ factor on the regret bound. M-OMD-H $(B=400)$ uses a larger memory budget, thereby performing better. The experimental results do not completely coincide with G2. It is left for future work to study whether we can avoid allocating the memory budget.
Table 2 also reports the average running time. It is natural that M-OMD-H $(B=400)$ requires the longest running time, since it uses a larger memory budget. The running time of the other three algorithms is similar.
Next we analyze the value of $\mathcal{A}_{T,\kappa^{\ast}}$. Computing $\mathcal{A}_{T,\kappa^{\ast}}$ requires $O(T^{2})$ time. For simplicity, we will define a proxy of $\mathcal{A}_{T,\kappa^{\ast}}$. Let
$\hat{\mathcal{A}}_{T,i}:=\sum^{T}_{\tau=1}\left\|\nabla_{\tau,i}-\hat{\nabla}_{\tau,i}\right\|^{2}_{\mathcal{H}_{i}}\cdot\mathbb{I}_{\nabla_{\tau,i}\neq 0}.$
To be specific, we use $\min_{i\in[K]}\hat{\mathcal{A}}_{T,i}$ as a proxy of $\mathcal{A}_{T,\kappa^{\ast}}$. In fact, our regret bound depends on $\hat{\mathcal{A}}_{T,i}$ which satisfies $\hat{\mathcal{A}}_{T,i}=\tilde{O}(\mathcal{A}_{T,\kappa_{i}})$. To obtain a precise estimation of $\mathcal{A}_{T,\kappa_{i}}$, we again run M-OMD-H with $M=30$ and $B=400$.
Table 2 shows the results. We find that $\mathcal{A}_{T,\kappa^{\ast}}<T$ on all datasets. Especially, $\mathcal{A}_{T,\kappa^{\ast}}\ll T$ on the mushrooms, the phishing, and the w8a datasets. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
@@
Raker [Shen2019Random] 11.97 $\pm$ 0.93 400 1.57 - 9.13 $\pm$ 0.28 400 2.06 -
LKMBooks [Li2022Worst] 6.04 $\pm$ 0.62 400 1.64 - 12.81 $\pm$ 0.54 400 2.21 -
@@
Raker [Shen2019Random] 32.14 $\pm$ 0.43 400 3.25 - 23.47 $\pm$ 0.23 400 9.45 -
LKMBooks [Li2022Worst] 23.66 $\pm$ 0.51 400 1.31 - 21.29 $\pm$ 1.36 400 10.15 -
@@
Raker [Shen2019Random] 2.98 $\pm$ 0.00 400 11.62 - 29.62 $\pm$ 0.46 400 8.88 -
LKMBooks [Li2022Worst] 2.97 $\pm$ 0.00 400 23.52 - 29.37 $\pm$ 0.87 400 5.47 -
@@
Raker [Shen2019Random] 9.49 $\pm$ 0.08 400 23.54 - 11.90 $\pm$ 0.31 400 53.32 -
LKMBooks [Li2022Worst] 9.58 $\pm$ 0.01 400 18.85 - 12.57 $\pm$ 0.26 400 19.17 -
@@ 6.2.2 Logistic Loss Function
We first compare the AMR of all algorithms. Table 3 reports the results. As a whole, M-OMD-S performs best on most of datasets. On the phishing, the SUSY, and the cod-rna datasets, Raker performs better than M-OMD-S. We have found that the current regret analysis of Raker is not tight and can be improved by utilizing Assumption 3. Thus it is reasonable that Raker performs similar with M-OMD-S. Besides, M-OMD-S performs better than LKMBooks on all datasets. The experimental results coincide with the theoretical analysis that is, M-OMD-S enjoys a better regret bound than LKMBooks. The experimental results verify G2. It is natural that the running time of all algorithms is similar as they enjoy the same time complexity.
Finally, we analyze the value of $L^{\ast}$. For simplicity, we use the cumulative losses of M-OMD-S, $\hat{L}_{1:T}$, as a proxy of $L^{\ast}$. Note that $L^{\ast}\leq\hat{L}_{1:T}$. Table 3 shows that $L^{\ast}<T$ on all datasets. Besides, on the mushrooms dataset, phishing dataset, and w8a dataset, $L^{\ast}\ll T$. In the optimal hypothesis space, a small memory budget is enough to guarantee a small regret and thus learning is possible. The experimental results verify G1.
@@
Raker [Shen2019Random] 11.47 $\pm$ 0.75 400 1.64 - 9.07 $\pm$ 0.44 400 2.05 -
LKMBooks [Li2022Worst] 10.01 $\pm$ 1.44 400 2.12 - 15.86 $\pm$ 1.28 400 1.94 -
@@
Raker [Shen2019Random] 30.74 $\pm$ 1.11 400 3.56 - 23.13 $\pm$ 0.17 400 9.90 -
LKMBooks [Li2022Worst] 25.72 $\pm$ 0.64 400 1.30 - 22.38 $\pm$ 1.29 400 10.98 -
@@
Raker [Shen2019Random] 2.97 $\pm$ 0.00 400 12.06 - 28.62 $\pm$ 0.44 400 9.02 -
LKMBooks [Li2022Worst] 2.98 $\pm$ 0.01 400 6.10 - 30.72 $\pm$ 1.33 400 6.45 0
@@
Raker [Shen2019Random] 9.53 $\pm$ 0.03 400 24.94 - 12.27 $\pm$ 0.57 400 52.24 -
LKMBooks [Li2022Worst] 9.57 $\pm$ 0.00 400 16.95 - 13.91 $\pm$ 1.31 400 25.55 -